OpenStudy (anonymous):

-11 +e^5x=17 I'm stuck between two answers? A.) X= in 28/in 5 B.) X= In 28/5

OpenStudy (anonymous):

I think it is B

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im not really sure how to do it, but ill try

OpenStudy (anonymous):

are those your answers

OpenStudy (anonymous):

choices*

OpenStudy (anonymous):

x= .666666 bro

OpenStudy (anonymous):

ln(28)/5

OpenStudy (oaktree):

So we can start by adding 11 to both sides, giving us the equation \[e^{5x} = 28\]From here, we can take the natural logarithm of both sides. \[\ln(e^{5x})=\ln(28) \rightarrow 5x = \ln(28) \rightarrow x=\frac{ \ln 28 }{ 5 }\] That should be an answer choice.

OpenStudy (oaktree):

Guys, try not to just give an answer, but explain where it comes from and how to get it.

OpenStudy (danjs):

log function is the inverse of an exponential ln is base e, natural log

OpenStudy (anonymous):

I was thinking B just wanted to make sure, thank you. @OakTree

OpenStudy (oaktree):

No problem!